# Disentangling data distribution for Federated Learning

### v1

@@ 5.1 Experimental settings
Models & Datasets. We conduct experiments on two datasets: CIFAR100 has the 20 superclass and each superclass has 5 subclass (Krizhevsky et al., 2014), thus, total 100 subclass; DomainNet (Z. Wu et al., 2015) has the 345 superclass (label) and each superclass has 5 subclass (style), thus, total 1725 subclass. We adopt ResNet (LeCun et al., 1998) for conducting the classification task to distinguish the superclass222each client only know the label of the superclass but doesn’t know the label of subclass on CIFAR100 and DomainNet. Please see details in Appendix A.
Federated Learning Settings. We simulate a horizontal federated learning system with K = 5, 10, 20 clients in a stand-alone machine with 8 Tesla V100-SXM2 32 GB GPUs and 72 cores of Intel(R) Xeon(R) Gold 61xx CPUs. For DomainNet and CIFAR10, we regard each subclass follow one sub-distribution ($P_{i}$). For the disentangled extent, we choose the averaged entangled coefficient $s=\frac{2}{K(K-1)}\sum_{k_{1},k_{2}}d_{k_{1},k_{2}}$ over all clients as $0,0.003,0.057$. The detailed experimental hyper-parameters are listed in Appendix A.
Baseline Methods. Four existing methods i.e., FedAvg (McMahan et al., 2017): FedProx (T. Li et al., 2020), SCAFFOLD (Karimireddy et al., 2020), MOON (Q. Li et al., 2021) and the proposed method FedDistr are compared in terms of following metrics.
@@ 6 Discussion and Conclusion
With advancements in computational capabilities during the era of large language models (LLMs) (Kasneci et al., 2023), the time consumption for local training has significantly decreased. Consequently, our focus is on enhancing communication efficiency. Moreover, the communication efficiency is especially important particularly in Wide Area Network scenarios (McMahan et al., 2017; Palmieri and Pardi, 2010).
Whether transferring data distribution leaks privacy is also an intriguing problem. Some studies (Xiao and Devadas, 2023) regard the underlying data distribution as a non-sensitive attribute, assuming that sharing insights into the distribution of data across users does not compromise individual privacy. However, other work indicates that using generative models to estimate the distribution still poses a risk of privacy leakage (Y. Wu et al., 2021).
@@ A.1 Dataset & Model Architectures
Models & Datasets. We conduct experiments on two datasets: CIFAR100 Krizhevsky et al. (2014) and DomainNet Z. Wu et al. (2015). For CIFAR100, we select 10 out of 20 superclasses, and 2 out of 5 subclasses in each superclass. For DomainNet Z. Wu et al. (2015), we select 10 out of 345 superclasses (labels), and 3 out of 5 subclasses (domains) in each superclass. We adopt ResNet LeCun et al. (1998) for conducting the classification task to distinguish the superclass333each client only know the label of the superclass but doesn’t know the label of subclass on CIFAR100 and DomainNet.
@@
Baseline Methods. Four existing methods i.e., FedAvg McMahan et al. (2017): FedProx T. Li et al. (2020), SCAFFOLD Karimireddy et al. (2020), MOON Q. Li et al. (2021) and the proposed method FedDistr are compared in terms of following metrics.

### v2

@@ 5.1 Experimental settings
Models & Datasets. We conduct experiments on two datasets: CIFAR100 has the 20 superclass and each superclass has 5 subclass (Krizhevsky et al., 2014), thus, total 100 subclass; DomainNet (Wu et al., 2015) has the 345 superclass (label) and each superclass has 5 subclass (style), thus, total 1725 subclass. We adopt ResNet (LeCun et al., 1998) for conducting the classification task to distinguish the superclass222each client only know the label of the superclass but doesn’t know the label of subclass on CIFAR100 and DomainNet. Please see details in Appendix A.
Federated Learning Settings. We simulate a horizontal federated learning system with K = 5, 10, 20 clients in a stand-alone machine with 8 Tesla V100-SXM2 32 GB GPUs and 72 cores of Intel(R) Xeon(R) Gold 61xx CPUs. For DomainNet and CIFAR100, we regard each subclass follow one sub-distribution ($P_{i}$). For the disentangled extent, we choose the averaged entangled coefficient $s=\frac{2}{K(K-1)}\sum_{k_{1},k_{2}}d_{k_{1},k_{2}}$ over all clients as $0,0.003,0.057$. The detailed experimental hyper-parameters are listed in Appendix A.
Baseline Methods. Four existing methods i.e., FedAvg (McMahan et al., 2017): FedProx (Li et al., 2020b), SCAFFOLD (Karimireddy et al., 2020), MOON (Li et al., 2021) and the proposed method FedDistr are compared in terms of following metrics.
@@ 6 Discussion and Conclusion
With advancements in computational capabilities during the era of large language models (LLMs) (Kasneci et al., 2023), the time consumption for local training has significantly decreased. Consequently, our focus is on enhancing communication efficiency. Moreover, the communication efficiency is especially important particularly in Wide Area Network scenarios (McMahan et al., 2017; Palmieri & Pardi, 2010).
Whether transferring data distribution leaks privacy is also an intriguing problem. Some studies (Xiao & Devadas, 2023) regard the underlying data distribution as a non-sensitive attribute, assuming that sharing insights into the distribution of data across users does not compromise individual privacy. However, other work indicates that using generative models to estimate the distribution still poses a risk of privacy leakage (Wu et al., 2021).
@@ A.1 Dataset & Model Architectures
Models & Datasets. We conduct experiments on two datasets: CIFAR100 Krizhevsky et al. (2014) and DomainNet Wu et al. (2015). For CIFAR100, we select 10 out of 20 superclasses, and 2 out of 5 subclasses in each superclass. For DomainNet Wu et al. (2015), we select 10 out of 345 superclasses (labels), and 3 out of 5 subclasses (domains) in each superclass. We adopt ResNet LeCun et al. (1998) for conducting the classification task to distinguish the superclass333each client only know the label of the superclass but doesn’t know the label of subclass on CIFAR100 and DomainNet.
@@
Baseline Methods. Four existing methods i.e., FedAvg McMahan et al. (2017): FedProx Li et al. (2020b), SCAFFOLD Karimireddy et al. (2020), MOON Li et al. (2021) and the proposed method FedDistr are compared in terms of following metrics.